### v1

@@ II.3 Operations modes
The nature of our quantum heat engine is dictated by the sign of the following quantities $Q_{c}$, $Q_{h}$, and $W$. As illustrated in Fig.(2), our engine can operate in one of the following four modes:
  • E
    (Heat engine): $Q_{c}\geq 0,Q_{h}\leq 0,W\geq 0$, the engine produces work by absorbing heat from the hot reservoir.
  • R
    ( Refrigerator): $Q_{c}\leq 0,Q_{h}\geq 0,W\leq 0$, the engine removes heat from the cold reservoir and expels it to the hot reservoir by converting the absorbed work.
  • A
    (Accelerator): $Q_{c}\geq 0,Q_{h}\leq 0,W\leq 0$, the engine absorbs energy and heats the cold reservoir from the hot reservoir.
  • H
    (Heater): $Q_{c}\leq 0,Q_{h}\leq 0,W\leq 0$, the engine absorbs heat and heats both reservoirs.
Figure 3: The different operating regimes of the quantum Otto engine in the $\alpha,\beta_{h}/\beta_{c},N-h_{i}$ plane. The color code represents the following modes: accelerator (green), heater (purple), heat engine (red), and refrigerator (blue).
@@
The purpose of this section is to show the switching between operational modes and their dependencies on the number of particles $N$, the interaction strength $\alpha$ , and the ratio $\beta_{h}/\beta_{c}$. All $\alpha$ values can be considered as long-range correlations, but for comparison purposes, we will divide them into three categories: long-range, intermediate, and short-range. This is because, as we increase $\alpha$, the effect of long-range correlations decreases and weakens. We have chosen specific values of $\alpha$, $\beta_{h}$, and $\beta_{c}$ where a diversity of modes can be observed in the ideal case. Considering the specific values imposed on our cycle, we observe the following behaviors:
For Fig.3($a$), where the system size is fixed at $N=100$, with $\beta_{h}=1.76$, $\beta_{c}=2.65$, and where $\alpha$ varies from 0 to 0.6, we observe different operational modes of the quantum engine as a function of the initial magnetic field $h_{i}$. For low values of $\alpha$ (between 0 and 0.2), long-range correlations between particles impose a refrigerator mode (blue) before the critical field $h=1$. For intermediate fields ($h_{i}=1.2$ to $1.7$), the accelerator (green) and heater (purple) modes appear; we see that they shift near the critical point to higher fields, playing the role of transient modes for the engine mode (red), which is predominant at high fields ($h_{i}=1.5$ to $2.5$). For intermediate $\alpha$ values (between $0.2$ and $0.4$), we observe a significant decrease in the refrigerator mode, indicating that medium-range correlations favor the appearance of the engine mode on both sides of the critical point ($h_{i}=0.5$ to $1.5$). The transient modes shift from $h_{i}=1.2$ at $\alpha=0.2$ to $h_{i}=0.2$ at $\alpha=0.4$. At intermediate fields, the heater and accelerator modes dominate, extending to high fields, with a significant decrease in the engine mode and the emergence of a refrigerator mode at high fields. For short-range correlations ($\alpha$ between $0.4$ and $0.6$), there is a slight decrease in the refrigerator mode and the disappearance of the transient modes A (accelerator) and H (heater), while the engine mode increases as $\alpha$ goes from $0.4$ to $0.6$. The H, A, and R (refrigerator) modes shift toward higher fields. The figure illustrates the impact of the magnetic field and correlation range on the operational modes.
In Fig.3($b$) the operational behavior is studied as a function of the ratio $\beta_{h}/\beta_{c}$, ranging from $0.45$ to $0.75$, with $\beta_{c}$ varying between 2.345 and 3.911, for a system size of $N=100$, with $\beta_{h}=1.76$ and $\alpha=0.25$. For higher values of $\beta_{h}/\beta_{c}$ (between $0.65$ and $0.75$), the refrigerator mode (blue) dominates and extends around the critical field at $h_{i}\approx 1.0$. The accelerator (A) and heater (H) modes also appear, acting as transitional modes between lower fields and the heat engine (E) in regions of more intense fields. As the $\beta_{h}/\beta_{c}$ ratio decreases (between 0.65 and 0.55), the refrigerator mode shows a significant reduction, while the A and H modes expand around the critical field, taking over from the high-field regions. Meanwhile, the zone corresponding to the engine mode (E) grows in the high-field regions and, to a lesser extent, in pre-transition phase chains. When the $\beta_{h}/\beta_{c}$ ratio is particularly low (between 0.45 and 0.55), the refrigerator mode rapidly declines for low fields, while the heater mode almost completely disappears. The accelerator mode, on the other hand, emerges after the phase transition, around $h_{i}\approx 1.5$, while the engine mode (E) strengthens, covering both low and high fields. These observations indicate that as the $\beta_{h}/\beta_{c}$ ratio approaches 1, the system tends to operate in refrigerator mode. However, as this ratio decreases, there is a gradual transition to heat engine mode, marked by the appearance and reinforcement of the accelerator and heater modes.
Fig.3($c$) examines the operational zones as a function of the system size $N$, which varies from $10$ to $100$, with $\alpha=0.25$, $\beta_{h}=1.76$, and $\beta_{c}=2.65$. For small system sizes (from $10$ to $40$), the refrigerator mode (blue) shows a dominance of the engine mode (red) for low fields before the critical point, with the appearance of the refrigerator mode at $N=20$ for low $h_{i}$. For intermediate fields, the accelerator (green) and heater (purple) modes appear (from $h_{i}=1.5$ to $2$), and the system operates exclusively in the refrigerator mode for strong fields. As we move to medium-length chains (from $40$ to $70$), there is a slight increase in the refrigerator mode for low $h_{i}$, with the appearance of the accelerator and heater modes and a slight retreat of the engine mode around the critical field. We also observe a continuous migration of the accelerator and heater modes from pre-critical fields toward the critical field, favoring the refrigerator mode, which appears in the intermediate and strong fields. For long chains (from $70$ to $100$), there is an increase in the refrigerator mode, which extends from low $h_{i}$ up to the critical field, with the disappearance of the engine mode after $N=92$ for low to intermediate fields. Additionally, the heater and accelerator modes expand from intermediate to strong fields, while the refrigerator mode barely appears and disappears in strong magnetic fields, favoring the engine, heater, and accelerator modes. We observe that fixing the range of interactions and the reservoir temperatures does not stabilize the operational modes since they also depend on the chain size.
This analysis reveals the intricate interplay between system size, interaction strength, and temperature ratio $\beta_{h}/\beta_{c}$ in determining the operational modes of a quantum engine. The results demonstrate that the operational modes are influenced not only by the chosen values for $\alpha$, $\beta_{h}$, and $\beta_{c}$, but also by the system size $N$. Through the analysis of the figures, it is evident that long-range correlations favor the refrigerator mode at low values of $\alpha$, while at intermediate values, the accelerator and heater modes become predominant. Additionally, the transient modes vary depending on the applied magnetic fields, illustrating complexdynamics when crossing the critical point. It has also been shown that varying the ratio $\beta_{h}/\beta_{c}$ significantly impacts the operational modes, favoring the refrigerator mode at higher values and leading to a transition to the heat engine mode at lower values. These results underscore the importance of a thorough understanding of interactions and correlations within the system to optimize the performance of quantum engines. This study paves the way for further explorations in optimization of the parameters to effectively exploit the different operational modes for practical applications in the field of quantum energy.
@@ II.4 Long Range Impact
Fig.4($a$) presents an analysis of the operational modes of a quantum thermal engine and a refrigerator in the ideal case. Initially, with $N$ fixed at $100$ and $\alpha$ varied from $0.01$ to $1.3$, we investigate the influence of interaction range and correlations on the efficiency and work output per particle $W/N$ of the heat engine, and on the efficacy and extracted heat per particle $Q_{c}/N$ of the refrigerator. Subsequently, we analyze the efficiency ($\eta$, $\eta_{R}$) as a function of $W/N$ and $Q_{c}/N$, respectively, for both the heat engine and refrigerator, considering system sizes $N$ ranging from $10$ to $100$. We analyze the different operational modes of a quantum heat engine and refrigerator in the ideal case; For the refrigerator, we consider $\beta_{h}=1.4$ and $\beta_{c}=1.5$ ($\beta_{h}=0.58$ and $\beta_{c}=0.6$), while for the thermal machines, $\beta_{h}=0.2$ and $\beta_{c}=0.9$ ($\beta_{h}=0.2$ and $\beta_{c}=0.667$).
Figure 4: The influence of long-range interactions on the work and the quantum refrigerator’s performance per spin as a function of $h_{i}$. Here $N$ is fixed at 100, while $\alpha$ is varied from 00 to $1.3$.
Figure 5: In the figure, $N$ is fixed at 100, and $\alpha$ is varied from 0 to 1.3 to observe the role of long-range interactions on the efficiency $\eta$ of the quantum heat engine and the coefficient of performance $\eta_{R}$ of the refrigerator as a function of $h_{i}$.
The ideal thermal machine is presented in the two figures, Fig.4($a$-$b$) and Fig.5($a$-$b$), where $W/N$ and $W/Q_{h}$ are again plotted as a function of $h_{i}$ for different ranges of $\alpha$. In this idealized scenario, we observe that for reservoirs with a large difference in reservoir temperatures (with $\beta_{h}=0.2$ and $\beta_{c}=0.667$), higher values of $\alpha$, which favor shorter correlations, result in higher work per spin $W/N$ compared to smaller and intermediate $\alpha$ before the critical point ($0.66$ for $\alpha=1.0$ to $0.77$ for $\alpha=1.3$).
There is a significant increase in $W/N$ for larger $h_{i}$, and near the critical field, a saturation of $W/N$ values occurs for larger $\alpha$, showing similar behavior. In the post-critical field region, an inversion occurs, with lower $\alpha$ ranges showing better $W/N$ than intermediate and higher ranges. This indicates that post-critical fields tend to favor dynamics that are more advantageous for short-range correlations. When the thermal gap is reduced (with $\beta_{h}=0.2$ and $\beta_{c}=0.9$), we observe higher work per spin for long-range correlations at very low magnetic fields, with a transition occurring before the phase transition, for short-range correlations, the work per spin $W/N$ improves, with a peak observed around $h_{i}=1$ at phase transition for long range correlations. The efficiency $\eta$ exhibits similar behavior, independent of the thermal gap. It is higher for long-range correlations, with a pronounced peak at low $h_{i}$. The thermal gap slightly shifts the peak but affects all correlations in the same way (see Fig.5($a$-$b$)). Let us now move on to the same study for the refrigerator in the case where the reservoires temperature difference is shorter, with $\beta_{h}=1.4$ and $\beta_{c}=1.5$, a similar pattern is observed for $\frac{Q_{c}}{N}$ , with greater work per spin for long-range correlations. The distinction becomes more evident in the efficiency $\eta_{R}$, which shows a marked rise for all $\alpha$ values, especially the higher ones. A notable saturation occurs at the critical point, causing a disruption and a decline in short-range correlations at $h_{i}=1.5$, rendering them less effective than long-range ones. At the same time, intermediate $\alpha$ values stabilize, experiencing a slight decrease but still outperforming the higher and short $\alpha$ values (red, blue).
In the case of the ideal refrigerator, Fig.4($c$-$d$) and Fig.5($c$-$d$), the first two figures depict the behavior where the ratio $\eta_{R}=\frac{Q_{c}}{|W|}$ and the coefficient of performance $\frac{Q_{c}}{N}$ are plotted against $h_{i}$ for different values of $\alpha$, with two very close reservoirs ($\beta_{h}=0.58$ and $\beta_{c}=0.6$). We observe that for low $h_{i}$ values (before $h_{i}=0.3$), short- and intermediate-range correlations are favored, allowing more heat to be extracted from the cold reservoir per spin compared to smaller $\alpha$. For slightly higher magnetic fields (from $h_{i}=0.3$ to $h_{i}=1$), there is a shift, and long-range correlations (small $\alpha$) begin to dominate. These long-range correlations saturate particularly near the critical field. After the phase transition, a significant increase in $\eta_{R}$ and $\frac{Q_{c}}{N}$ is observed for all $\alpha$ values, without a reversal. Starting from $h_{i}=1.5$, stabilization occurs, with a slight decrease in performance for large and intermediate $\alpha$, while small $\alpha$ maintains its effectiveness, indicating that long-range correlations are more favorable in the post-critical region.
In an ideal refrigerator, we also observe that for small values of $\alpha$ (e.g., from 0.1 to 0.4), both $\frac{Q_{c}}{|W|}$ and $\frac{Q_{c}}{N}$ gradually increase with $h_{i}$, reaching a peak for higher values of $\alpha$. As $\alpha$ increases, the system is able to maintain its cooling efficiency over a wider range of $h_{i}$. This suggests that, under ideal conditions, the refrigerator can sustain a stable cooling effect even as the range of interactions extends, with improved performance observed for higher values of $\alpha$.
Figure 6: Efficient $\eta$ (COP $\eta_{R}$) as a function of $h_{i}$ and W/N ($Q_{c}$/N).
In Fig.6($c$-$d$), which illustrates the effective efficiency $\eta_{R}$ as a function of $Q_{c}/N$ and the initial transverse field $h_{i}$, small values of $N$ (around 10-30) exhibit a high sensitivity to variations in $\frac{Q_{c}}{N}$ and $h_{i}$. The effective efficiency $\eta_{R}$ reaches elevated values within a narrow range of $Q_{c}/N$, indicating an unstable behavior of the system for these smaller sizes. As $N$ increases, the curves become smoother, suggesting a stabilization of the efficiency and a reduction in sensitivity to fluctuations. For higher values of $N$, the maximum value of $\eta_{R}$ shifts towards lower $Q_{c}/N$, indicating that optimal efficiency is achieved with lower heat extraction per cycle, likely due to a more effective management of interactions in a larger system. We observe a sequential behavior: before the phase transition, we have higher efficiency $\eta_{R}$ for short chains and larger heat extraction $Q_{c}/N$ from the cold reservoir for long chains. At the critical field $h_{i}=1$, a shift occurs, and the behavior changes. Long chains become more efficient, with $\eta_{R}$ exceeding that of short chains, and $Q_{c}/N$ becomes significant for small $N$, indicating an $N$-dependence of the efficiency after the phase transition. For the second case with $\alpha=1.2$, we observe a distinct behavior and greater stability compared to the previous case. This indicates that as $\alpha$ increases, short-range correlations are favored, making them less sensitive to thermal fluctuations. Consequently, they become less dependent on the phase transition, which primarily enhances efficiency and $Q_{c}/N$, both of which are more significant for short chains.
In Fig.7($a$-$b$) In the following, we will comment on the figure, which represents the thermodynamic efficiency $\eta$ plotted as a function of $h_{i}$ and the work per spin $\frac{W}{N}$. For small values of $N$ ($10$ to $40$), at low $h_{i}$, we observe a significant efficiency, although lower than that of intermediate and longer chains, with a higher work per spin for smaller $N$. As the magnetic field increases, there is a sharp drop for short chains near the critical field $h=1$, with a slight decrease for larger $N$, which increases fluctuations for small $N$ and quickly lowers $\frac{W}{N}$ for stronger chains. Meanwhile, for longer and intermediate chains, at the critical field, there is a drop in efficiency and a significant increase in the work per spin $\frac{W}{N}$. This indicates that the efficiency of the cycle strongly depends on the initial applied field and the size of the chain.
In this study, we have explored the operational modes of a quantum heat engine and a refrigerator, emphasizing the effects of long- and short-range correlations as governed by the parameter $\alpha$. By examining various values of $\alpha$ and different reservoir temperature configurations, we have identified distinct behaviors in both the quantum thermal machine and the refrigerator.
For the quantum heat engine, long-range correlations (small $\alpha$) enhance performance in the pre-critical region, while short-range correlations (larger $\alpha$) dominate in the post-critical region. This demonstrates that critical dynamics significantly impact the work per spin $W/N$ and the efficiency. In the case of larger reservoir temperature differences, $\alpha$ plays a crucial role in determining performance, especially near the critical point where saturation is observed.
For the refrigerator, long-range correlations are similarly favored post-critical field, particularly in terms of cooling capacity $Q_{c}/N$ and efficiency $\eta_{R}$. The proximity of the reservoirs’ temperatures reveals that long-range correlations maintain stable performance even after the critical transition, while intermediate and short-range correlations exhibit a decline in efficiency.
Overall, the analysis highlights the significant influence of correlation range and magnetic field strength $h_{i}$ on the thermodynamic properties of both systems. Long-range interactions provide distinct advantages in certain regimes. Moreover, the stabilization of these effects for larger system sizes $N$ suggests that larger chains offer better optimization of quantum thermal devices, while smaller chains are more sensitive to fluctuations, particularly near the critical field. This study provides valuable insights into optimizing quantum heat engines and refrigerators through the control of system parameters and interactions.

### v2

@@ II.3 Operations modes
The purpose of this section is to show the switching between operational modes and their dependencies on the number of particles $N$, the interaction strength $\alpha$ , and the ratio $\beta_{h}/\beta_{c}$.Let us now analyze the case of an infinitely slow cycle , the limit $t\to\infty$. All $\alpha$ values can be considered as long-range correlations, but for comparison purposes, we will divide them into three categories: long-range, intermediate, and short-range. The color code represents the following modes: accelerator (green), heater (purple), heat engine (red), and refrigerator (blue).As illustrated in Fig.3, our engine can operate in one of the following four modes:
  • Refrigerator (R): the engine absorbs energy and transfers heat from the cold reservoir to the hot one, i.e., $Q_{c}>0$, $Q_{h}<0$, and $W<0$;
  • Accelerator (A): the engine absorbs energy and transfers heat from the hot reservoir to the cold one, i.e., $Q_{c}<0$, $Q_{h}>0$, and $W<0$;
  • Heat engine (E): the engine produces work by absorbing heat from the hot reservoir, i.e., $Q_{c}<0$, $Q_{h}>0$, and $W>0$;
  • Heater (H): the engine absorbs energy and heats up both the hot and the cold reservoirs, i.e., $Q_{c}<0$, $Q_{h}<0$,and $W<0$.
Figure 4-a summarises the machine’s operational regimes for $N=100$, $T_{h}=0.57$ and $T_{c}=0.38$ as the interaction range is varied. Physically, decreasing $\alpha$ (making the interactions longer-ranged) strengthens collective correlations and coherent energy exchange across the chain: this favors refrigeration at low fields and promotes nonlocal response modes at intermediate fields, because the system can coordinate energy flows across many sites. By contrast, increasing $\alpha$ (shortening the interaction range) suppresses these collective channels, which reduces the region where refrigeration and transient modes appear and instead stabilises steady engine-like operation over a wider field window. In other words, longer-range couplings enhance sensitivity to global field conditions and produce pronounced finite-$N$ modes, whereas more local couplings favour robust, local energy extraction and smoother, more extensive engine performance. The former case $\alpha=0.2$, lies well outside the perturbative domain of Fisher’s expansion and corresponds to an interaction range so extended that non-additive, quasi–mean-field behaviour is expected; the latter, $\alpha=1.2$, remains within the classical long-range sector but with markedly weaker long-range effects. These conclusions are therefore fully compatible with the analytical expectations of [10] and with the qualitative differences observed in our numerical data. Figure 4-b displays the machine’s operational map for $N=100$ and $\alpha=0.25$, with $T_{h}=0.57$ while the cold temperature is varied so that the ratio $T_{c}/T_{h}$ spans 0.450.450.750.75. Physically, changing $T_{c}/T_{h}$ tunes the thermodynamic bias available for work extraction: larger ratios (temperatures closer together) reduce the net thermal gradient and suppress steady work output, making thermally-driven refrigeration and transient response modes more prominent near the critical field. Conversely, decreasing $T_{c}/T_{h}$ (increasing the thermal contrast) strengthens the directional heat flow that drives net work production, stabilising engine-like operation across a broader field range and diminishing the parameter window where refrigerator or heater modes appear. Transient accelerator/heater bands that flank the critical region reflect a competition between critical susceptibility (which amplifies small driving imbalances) and the global thermal bias; as the thermal bias grows, this competition is resolved in favour of persistent engine behaviour. Figure 4-c shows how the machines operating regimes reorganize as the system size $N$ is varied (here $\alpha=0.25$, $T_{h}=0.57$, $T_{c}=0.38$). Physically, the observed changes reflect a competition between two size-dependent effects: (i) finite-$N$ coherence and level discreteness, which favour directed, coherent energy extraction when the chain is short, and (ii) the progressive densification of the many-body spectrum and the growth of low-energy collective modes as $N$ increases, which amplify critical susceptibility and thermal redistribution near the critical field. For small chains, the sparsity of excitations and enhanced coherence make steady engine-like operation energetically favourable at low fields. As $N$ grows the spectrum becomes denser and low-energy fluctuations proliferate; these modes couple more efficiently to the reservoirs and to field-induced critical responses, expanding refrigeration and transient accelerator/heater bands close to criticality and reducing the parameter window for steady engine operation. At the largest sizes the increased phase-space for scattering and thermal equilibration further reinforces refrigeration-dominated behaviour over a broad field interval. In short, increasing $N$ shifts the device from a regime dominated by finite-size coherent extraction to one governed by many-body critical fluctuations and thermal redistribution, which explains the progressive replacement of engine by refrigerator behaviour in the data.The device’s operational mode is set by three parameters: the interaction exponent $\alpha$, the thermal contrast $T_{c}/T_{h}$ (or $\Delta T$), and the system size $N$. Smaller $\alpha$ (long-range couplings) enhances global coherence and favours refrigeration with pronounced finite-$N$ transients near criticality, whereas larger $\alpha$ suppresses collective channels and stabilises engine-like operation. A small thermal gap (high $T_{c}/T_{h}$) weakens net work extraction and biases the system toward refrigeration or transient responses, while a large gap strengthens sustained engine performance. Small $N$ permits coherent finite-size windows for directed extraction; large $N$ increases low-energy fluctuations and thermalisation phase-space, broadening refrigeration and transient regimes. Hence $\alpha$, $T_{c}/T_{h}$, and $N$ constitute direct, tunable levers for steering and optimising quantum thermal machines.
Figure 7: Efficient $\eta$ (COP $\eta_{R}$ ) as a function of $h_{i}$ and W/N ($Q_{c}$/N) this highlights the behavior of the efficiency for different initial field values $h_{i}$ and of the work per spin $W/N$, as well as the possible optimization paths of the efficiency for various interaction ranges characterized by different values of $\alpha$. .
@@ II.4 Long Range Impact
In the present study, we explore how the distinctive attributes of long-range interacting quantum systems can enhance the efficiency of a many-body quantum Otto cycle, adopted here as a representative model of a quantum thermal device. Our analysis reveals several key benefits similar to [9], derived from the incorporation of long-range interactions:
  1. 1.
    The infinitely slow limit , the device attains markedly improved optimal performance. This enhancement is most evident in the practically relevant heat-engine and refrigerator regimes when collective and local effects are optimally balanced. .
  2. 2.
    The interaction-range exponent $\alpha$ affects the cycle efficiency differently depending on the operational mode.
Relying on the results reported by [11], it was shown that the spatial decay of connected correlations depends sensitively on $\alpha$ and on proximity to the critical phase transition: large $\alpha$ yields short-range exponential screening, small $\alpha$ produces algebraic long-range tails, and an intermediate hybrid crossover combines both behaviours. On the paramagnetic side, quasi-exponential screening amplifies nearest-neighbour influence, whereas sufficiently small $\alpha$ restores system-wide coherence and collective channels. Crucially, the presence of a spectral gap does not preclude algebraic tails, so that spectral structure intrinsically limits the efficacy of nonlocal propagation. In Figures 5 and 6, thermodynamic performance is plotted as a function of the initial magnetic field $h_{i}$ for various values of $\alpha$. When the temperature gap is large ($T_{h}=5.0$, $T_{c}=1.11$) , an increase in $\alpha$, corresponding to a reduction in the range of correlations, leads to a higher work output per spin before the critical region. At the critical phase, an inversion occurs where long-range interactions with smaller $\alpha$ outperform those with larger $\alpha$, suggesting that post-critical dynamics favor long-range interactions. This is explained by the influence of neighboring particles, which is more effective in the paramagnetic phase, where increased spin misalignment enhances work extraction. When the thermal gap is reduced (e.g., $T_{c}=0.5$, $T_{h}=1.67$), long-range interactions yield a higher $W/N$, as an increase in $\alpha$ induces localized interactions that propagate more weakly than long-range interactions when the thermal gap is small. The efficiency $\eta$ exhibits a similar trend across configurations, reaching a maximum for long-range correlations at low $h_{i}$, with a slight shift in the peak depending on the thermal gap. Finally, in the refrigerator configuration with a narrow thermal gap ($T_{h}=0.71$, $T_{c}=0.67$ and $T_{h}=1.72$, $T_{c}=1.67$), a similar behavior is observed for the cooling power $Q_{c}/N$, with long-range interactions yielding superior performance, as the extended range of interactions facilitates more effective heat diffusion compared to larger $\alpha$ values, which prioritize localized transmission. An increase in $\alpha$ for the efficiency $\eta_{R}$, which reduces dissipation to nearest neighbors, significantly enhances the yield until it reaches a saturation threshold. Guided by the conclusions of [12, 13], the observed saturation of the critical peaks is a direct consequence of finite-temperature rounding: the thermal length scale $\xi_{T}$ bounds the spatial correlations, so that finite-size scaling applies only for $N\lesssim\xi_{T}$, whereas for $N\gg\xi_{T}$ would-be critical divergences are suppressed and instead manifest as extended plateaus in the relevant observables. Long-range couplings (small $\alpha$) thus enhance performance for short chains but produce rapid saturation once $N$ exceeds $\xi_{T}$; by contrast, larger $\alpha$ yields a more gradual increase in performance that ultimately levels off. Namely the coexistence in the QLRO phase of violations of the area law for entanglement entropy and a gapped entanglement spectrum are consistent with a picture in which the spread of quantum correlations is limited. This limited spread explains the saturation of the performance peaks we observe and suggests the existence of an energetic barrier that naturally limits the spread of quantum correlations.
Figure 8: Contour plot of $Q_{c}$ denotes the the heat exchanged during the thermalisation with the cold bath done by the system and $\eta_{R}$ as a function of $h_{i}$ for different values of $\alpha$, with (a) $\alpha=0.25$, (b) $\alpha=1.2$, (c) $\alpha=0.25$, and (d) $\alpha=1.2$.
Figure 9: COP of the refrigerator versus $h_{i}$, for different system sizes NN across the critical point. The critical field is located at $h_{i}=1$ for $\alpha=0.25$, and approximately at $h_{i}\approx 0.75$ for $\alpha=1.2$.
To better understand optimization pathways, the behavior of thermodynamic efficiency $\eta$ and work per spin $W/N$ is analyzed as a function of the transverse field $h_{i}$ for different $N$ in Fig. 7.a-b. For small $N$, high values of $\eta$ are observed at low $h_{i}$, despite a higher $W/N$ compared to larger systems. Near the critical point, $h_{i}=1$, short chains experience a sharp drop in efficiency, whereas longer chains exhibit a more gradual transition with a marked increase in $W/N$, displaying a more linear dependence on $N$. This contrasts with the refrigerator mode in Fig. 7.c-d, where short chains yield a higher $\eta_{R}$, while longer chains extract more heat. Post-phase transition, longer chains predominate due to dominant short-range correlations, which reduce sensitivity to phase transitions and thermal fluctuations. In summary, this work has investigated the performance of quantum heat engines and refrigerators by analyzing the role of interaction range—encoded in the parameter $\alpha$—across different thermodynamic regimes.